An electric current of 2.0 A passes through a wire of resistance 25 Ω. How much heat will b developed in 1 minute?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. i = 2 A, r = 25 Ω, T = 1 min = 60 sec Heat developed = i^2 RT = 2 * 2 * 25 * 60 = 6000 J